A 1000 kg automobile enters a freeway on ramp at 20 m/s and accelerates uniformly up to 40 m/s in a time of 10 sec. How far does the automobile travel during this time ?
1
Expert's answer
2013-01-11T06:57:41-0500
Due to the law of the point movement:X-coordinate
Xo- initial coordinate
Vo - initial velocity
t - time of moving
a - acceleration
m - mass - isn't necessary for this task
X=Xo+Vo*t+a*t^2/2 - law of the point movement But we need to find the way travelled, so we assume Xo=0 Then X=Vo*t+a*t^2/2=20*10+40*100/2=220m